# Flatness-based Finite-Horizon Multi-UAV Formation Trajectory Planning and Directionally Aware Collision Avoidance Tracking

### v1

@@ 7 Conclusion
In this paper, we present a collision-free finite-time formation control scheme for multiple UAVs based on the differential flatness property of UAV dynamics. Unlike traditional approaches that rely on a leader, predefined tracking signals, or waypoints, our method leverages the initial states of the UAVs to plan optimal formation trajectories. Moreover, the proposed approach introduces a directionally aware collision avoidance strategy that prioritizes collision avoidance both in the forward path and the relative approach for each UAV, promoting adaptive behavior and enhancing efficiency. By addressing a finite-time optimal control problem, we derived a trajectory that optimizes a collective performance index, ensuring efficient formation and movement. Additionally, we tackled the collision-constrained regulating problem to maintain collision-free tracking of the formation trajectory. We trade optimality for safety guarantees in the tracking controller when necessary. The simulation results validated the efficacy of our control scheme.
The assumption of a complete communication graph to ensure global collision avoidance awareness may result in scalability challenges for larger UAV teams. Future work could address this limitation by developing distributed control strategies that enable real-time, directionally aware, collision-free trajectory planning and tracking for physical UAVs using onboard sensors such as UVDAR [27].
@@ Acknowledgments
$\mathbf{r}(t)=\begin{bmatrix}\mathbf{I}&\mathbf{0}\end{bmatrix}\exp{((t-t_{f})\mathbf{M})}\begin{bmatrix}\mathbf{I}\\ \mathbf{Q}_{f}\end{bmatrix}\mathbf{H}^{-1}(t_{f})\mathbf{r}_{0}.$ (50)
Using the notation in (9), we get (8). This concludes the proof. ∎

### v2

@@ 7 Conclusion
This paper presents a collision-free, finite-time formation control scheme for multiple UAVs based on the differential flatness property of UAV dynamics. Unlike traditional methods that rely on a leader, predefined tracking signals, or waypoints, our approach utilizes the initial states of the UAVs to plan optimal formation trajectories based on consensus. Additionally, we introduce a directionally aware collision avoidance strategy that prioritizes avoiding collisions both in the forward path and the relative approach for each UAV, promoting adaptive behavior and enhancing efficiency. By formulating and solving a finite-time optimal control problem, we derive trajectories that optimize a collective performance index, ensuring efficient formation and movement of the UAV team. To maintain collision-free tracking of the formation trajectory, we incorporate safety guarantees into the tracking controller when necessary, striking a balance between optimality and safety. High-fidelity simulations are performed using the ROS-Gazebo-based simulator. The results confirm the effectiveness of our control scheme, demonstrating successful formation trajectory planning and formation trajectory tracking while avoiding collisions. However, the assumption of a complete communication graph for global collision awareness may pose scalability challenges for larger UAV teams. Future work could address this limitation by developing distributed control strategies that enable real-time, directionally aware, collision-free trajectory planning and tracking, using onboard sensors such as UVDAR [41].
@@ Acknowledgments
$\mathbf{r}(t)=\begin{bmatrix}\mathbf{I}&\mathbf{0}\end{bmatrix}\exp{((t-t_{f})\mathbf{M})}\begin{bmatrix}\mathbf{I}\\ \mathbf{Q}_{f}\end{bmatrix}\mathbf{H}^{-1}(t_{f})\mathbf{r}_{0},$ (52)
$\bm{\psi}(t)=\begin{bmatrix}\mathbf{0}&\mathbf{I}\end{bmatrix}\exp{((t-t_{f})\mathbf{M})}\begin{bmatrix}\mathbf{I}\\ \mathbf{Q}_{f}\end{bmatrix}\mathbf{H}^{-1}(t_{f})\mathbf{r}_{0}.$ (53)
Using the notation in (10) and (11), we get (8) and (9). This concludes the proof. ∎
Appendix B stability analysis of forward-path-aware collision avoidance control
This appendix analyzes the stability of the collision avoidance control strategy under the forward-path-aware penalty function $v_{ij}^{F}$, as introduced in Subsection 5.1. Specifically, it demonstrates that the control inputs remain bounded without relying on Assumption 4, ensuring the stability of the directionally aware collision avoidance framework. The analysis focuses on the gradient of the penalty function and its impact on the control law in (23).
Recall the collision avoidance control strategy (23) under the forward-path-aware penalty function as follows,
$\mathbf{u}_{\mathbf{z}}(t)=-\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\mathbf{P}(t)\mathbf{z}(t)-\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\frac{\partial V^{F}}{\partial\mathbf{z}}^{\top},$ (54)
where $V_{i}^{F}=\sum_{j\in\mathcal{N}_{i}}v_{ij}^{F}=\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}v_{ij}$ for UAV $i$ and the total penalty is $V^{F}=\sum_{i}V_{i}^{F}$.
Then, we have
$\frac{\partial V_{i}^{F}}{\partial\mathbf{z}_{i}}=\sum_{j\in\mathcal{N}_{i}}\big{(}\alpha_{ij}\underbrace{\frac{\partial v_{ij}}{\partial\mathbf{z}_{i}}}_{\text{original gradient}}+v_{ij}\underbrace{\frac{\partial\alpha_{ij}}{\partial\mathbf{z}_{i}}}_{\text{directional adjustment}}\big{)}.$ (55)
The gradient decomposition in (55) shows how directional awareness modifies the collision response. The original gradient term maintains baseline collision avoidance, while the directional adjustment term modulates responses based on the forward-path awareness.
To ensure stability and avoid unbounded control inputs, we must guarantee that the gradient of the total penalty $\frac{\partial V^{F}}{\partial\mathbf{z}}$ does not grow indefinitely.
The total penalty $V$ is guaranteed to be non-increasing because of bounded components $\alpha_{ij}$ and $v_{ij}$ (where $\alpha_{ij}\in[0,1]$ and $v_{ij}\to 0$ as $\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\to\infty$) and vanishing gradients
$\left\|\frac{\partial V_{i}^{F}}{\partial\mathbf{z}_{i}}\right\|\leq\sum_{j\in\mathcal{N}_{i}}\left(\alpha_{ij}\left\|\frac{\partial v_{ij}}{\partial\mathbf{z}_{i}}\right\|+v_{ij}\left\|\frac{\partial\alpha_{ij}}{\partial\mathbf{z}_{i}}\right\|\right),$
where $\frac{\partial v_{ij}}{\partial\mathbf{z}_{i}}$ is proven to be bounded in Lemma 1, and for $\frac{\partial\alpha_{ij}}{\partial\mathbf{z}_{i}}$, its bounds are derived as follows.
For $\alpha_{ij}=\frac{(\mathbf{p}_{j}-\mathbf{p}_{i})^{\top}\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|}$ (when $\|\mathbf{v}_{i}\|\neq 0$ and $\cos\eta_{ij}>0$), the partial position and velocity derivatives are
$\frac{\partial\alpha_{ij}}{\partial\mathbf{p}_{i}}=-\frac{\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|}+\frac{(\mathbf{p}_{j}-\mathbf{p}_{i})^{\top}\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|^{3}\|\mathbf{v}_{i}\|}(\mathbf{p}_{j}-\mathbf{p}_{i}),$
$\frac{\partial\alpha_{ij}}{\partial\mathbf{v}_{i}}=\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|}-\frac{(\mathbf{p}_{j}-\mathbf{p}_{i})^{\top}\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|^{3}}\mathbf{v}_{i}.$
Taking their Euclidean norm, we have
$\left\|\frac{\partial\alpha_{ij}}{\partial\mathbf{p}_{i}}\right\|$ $\leq\left\|-\frac{\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|}\right\|+\left\|\frac{(\mathbf{p}_{j}-\mathbf{p}_{i})^{\top}\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|^{3}\|\mathbf{v}_{i}\|}(\mathbf{p}_{j}-\mathbf{p}_{i})\right\|$
$=\frac{1}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|}+\frac{|\cos\eta_{ij}|}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|}\leq\frac{2}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|},$
and
$\left\|\frac{\partial\alpha_{ij}}{\partial\mathbf{v}_{i}}\right\|$ $\leq\left\|\frac{\mathbf{p}_{j}-\mathbf{p}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|}\right\|+\left\|\frac{(\mathbf{p}_{j}-\mathbf{p}_{i})^{\top}\mathbf{v}_{i}}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\|\mathbf{v}_{i}\|^{3}}\mathbf{v}_{i}\right\|$
$=\frac{1}{\|\mathbf{v}_{i}\|}+\frac{|\cos\eta_{ij}|}{\|\mathbf{v}_{i}\|}\leq\frac{2}{\|\mathbf{v}_{i}\|}.$
Thus, the total derivative norm is bounded by
$\left\|\frac{\partial\alpha_{ij}}{\partial\mathbf{z}_{i}}\right\|\leq\left\|\frac{\partial\alpha_{ij}}{\partial\mathbf{p}_{i}}\right\|+\left\|\frac{\partial\alpha_{ij}}{\partial\mathbf{v}_{i}}\right\|\leq\frac{2}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|}+\frac{2}{\|\mathbf{v}_{i}\|}.$
Since $v_{ij}\to 0$ as $\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\to\infty$, the term $v_{ij}\frac{\partial\alpha_{ij}}{\partial\mathbf{z}_{i}}$ in (55) diminishes at large separations, further ensuring that the directional adjustment term does not contribute to unbounded growth in the total gradient.
The boundedness of $\frac{\partial\alpha_{ij}}{\partial\mathbf{z}_{i}}$, combined with the boundedness of $\frac{\partial v_{ij}}{\partial\mathbf{z}_{i}}$ (as proven in Lemma 1) and the decay of $v_{ij}$ as $\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\to\infty$, ensures that the total gradient $\frac{\partial V_{i}^{F}}{\partial\mathbf{z}_{i}}$ in (55) remains bounded. Consequently, the control input $\mathbf{u}_{\mathbf{z}}(t)$ in (54) avoids unbounded growth, preserving the stability of the HJB equation and the collision avoidance strategy without Assumption 4.
Similar derivations can be applied to the approach-aware penalty function $v_{ij}^{A}=\beta_{ij}v_{ij}$ and the unified penalty function $v_{ij}^{U}=\xi_{ij}v_{ij}$, as their weighting factors ($\beta_{ij}$, $\xi_{ij}$) are also bounded in $[0,1]$ and their gradients can be shown to be bounded using analogous techniques, ensuring stability across the directionally aware framework.